OpenStudy (anonymous):

*I'll give medal and fan!* Solve: 11/13 - 4/7 . Answer must be a simplified fraction.

OpenStudy (anonymous):

\[\huge \frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bd}\]

OpenStudy (anonymous):

\[\huge \frac{11}{13}-\frac{4}{7}=\frac{11\times 7-13\times 4}{13\times 7}\]

OpenStudy (anonymous):

that is the one and only way to subtract fractions i will leave the rest of the calculation to you